### v1

@@ Abstract
We study how generative AI affects labor market signaling using the introduction of an AI-powered cover letter writing tool on Freelancer.com. Our data track both access to the tool and usage at the application level. Difference-in-differences estimates show that access to the AI tool increased textual alignment between cover letters and job posts—which we refer to as cover letter tailoring—and raised callback likelihoods. Workers with weaker pre-AI writing skills saw larger improvements in cover letters, indicating that AI substitutes for workers’ own skills. Although only a minority of applications used the tool, the overall correlation between cover letter tailoring and callbacks fell by 51%, implying that cover letters became less informative signals of worker ability in the age of AI. Employers correspondingly shifted toward alternative signals, such as workers’ past reviews, which became more predictive of hiring. Finally, within the treated group, greater time spent editing AI drafts was associated with higher hiring success.
@@ Introduction
Job seekers and college applicants increasingly turn to generative artificial intelligence (AI) tools like ChatGPT to edit résumés, craft cover letters, and prepare for interviews. [^3] This trend has received widespread media attention. Recent headlines include “AI is supposed to make applying to jobs easier—but it might be creating another problem” (NBC News, November 17, 2024) and “AI Is Taking Over College Admissions” (The Nation, October 11, 2024). In 2025, LinkedIn reported a 45 percent year-over-year increase in job applications, with generative AI widely believed to be a key driver of the surge (“Employers Are Buried in A.I.-Generated Résumés,” The New York Times, June 21, 2025). Despite the transformative potential of this technology for processes central to economic opportunities and social mobility, credible empirical evidence on how AI affects labor market signaling remains scarce. [^4] Notable exceptions include Cowgill et al. (2024), who provide evidence on applicants’ use of generative AI in a lab setting, and Wiles and Horton (2025), who study the effect of employers’ use of generative AI tools to write job posts on an similar online labor platform. The main obstacle is observability: outside of controlled lab settings, just as employers struggle to identify AI-assisted applications, so too do researchers. Current detection tools are error-prone and unreliable (Weber-Wulff et al. (2023)), making it difficult to assess AI assistance’s real-world impact on hiring.
@@
To measure how closely cover letters are tailored to job descriptions, we apply Term Frequency–Inverse Document Frequency (TF-IDF), a standard Natural Language Processing technique that captures the extent of keyword overlap between two texts. Our approach is related to, but distinct from, Galdin and Silbert (2024), who use data from the same platform and ask a Large Language Model (LLM) whether a given cover letter aligned with a job posting. They show that the LLM’s estimated probability of a “yes” response predicts hiring outcomes. [^6] They use jobs listed under the skill categories PHP, Python, and Data Entry, posted between January 1, 2019 and February 29, 2020—a period earlier than our data sample. Using data from before the introduction of the AI writing tool, we show that our TF-IDF–based measure of cover-letter tailoring similarly predicts callbacks and awards—both across and within workers—providing complementary evidence that employers value cover letters that engage with job content.
@@
We find that access to the generative AI writing tool increased cover-letter tailoring by 0.16 standard deviations, while actual usage raised tailoring by 1.36 standard deviations. Applying the same design to callbacks as the outcome, we find that access to the generative AI tool increased the probability of receiving a callback by 0.43 percentage points, and usage raised it by 3.56 percentage points. The latter represents a 51% increase relative to the pre-rollout average callback rate of 7.02%. However, the results for callbacks are significant at the $10\%$ level but not at the $5\%$ level. A dynamic effect specification shows that the effect on callbacks tapered off after two months. These results suggest that generic or poorly matched cover letters present a barrier to receiving callbacks. In contrast, the estimates on actual awards are too imprecise to draw definitive conclusions.
Our second finding is that AI substitutes for, rather than complements, workers’ pre-AI cover letter tailoring skills. We compute workers’ average tailoring scores prior to the tool’s introduction and study the heterogeneous effects of access to the tool by pre-existing writing ability. We find that workers who previously wrote more tailored cover letters experienced smaller gains in cover letter tailoring—indeed, the best writers gained experienced 27% smaller gains than the weakest ones. By enabling less skilled writers to produce more tailored cover letters, AI narrows the gap between workers with different initial abilities.
@@
Our paper builds on Galdin and Silbert (2024), who study how market congestion affects workers’ effort in customizing cover letters to signal their fit with jobs, using data from Freelancer. They are the first to measure alignment between cover letters and job posts in this context—employing a different method from our TF-IDF-based measure—and show that cover letter alignment predicts employers’ hiring decisions. To analyze the effect of market congestion, Galdin and Silbert (2024) use a structural signaling model and note that their framework can simulate counterfactual scenarios in which workers use AI assistance. In contrast, we leverage the real-world introduction of an on-platform AI writing tool to directly evaluate its impact on cover letter quality and effectiveness.
@@
A few recent papers study the effect of AI on labor matching. The most closely related work is Cowgill et al. (2024), who examine the impact of generative AI on job applications and startup pitches through a survey experiment. They find that generative AI reduces employers’ and investors’ screening accuracy. Our study explores similar questions in an online labor market, examining the effect of AI on actual hiring decisions. An additional important difference between the settings is that Cowgill et al. (2024) informed employers whether applicants used generative AI, whereas, in many real-world contexts—including ours—employers typically cannot discern whether an applicant received assistance from AI. As a result, our findings are not directly comparable.
@@
In a similar labor market, Wiles et al. (2023) study the effect of a non-generative algorithmic writing assistance on employment through an experiment. The intervention they study focuses mostly on correcting spelling and grammar errors. They show that the tool improved treated workers’ employment outcomes without affecting the employers’ satisfaction. While we also investigate the role of AI—although of a different nature—in workers’ employment outcomes, we are additionally able to speak to the market-level outcomes associated with the introduction of generative AI. Another related paper is Wiles and Horton (2025), who study generative AI assistance to employers for crafting job posts in a similar context and find that while the tool increased the number of job posts, it did not increase the number of matches. Our work is complementary to theirs as both sides of the labor market—firms and workers—increasingly adopt generative AI tools in matching.
@@
Also related is work that examines the functioning of online labor markets and their broader economic implications. Horton (2010), Chen and Horton (2016), and Stanton and Thomas (2020) document how these markets operate, the behavioral responses of workers to wage changes, and the significance of platform-mediated employment. Stanton and Thomas (2025) provide evidence on the distributional consequences of gig platforms, showing how access to online markets affects both workers and firms, and highlighting heterogeneity in who ultimately benefits from platform participation. This focus on heterogeneous impacts complements theoretical and empirical work such as Horton et al. (2024) and Allman et al. (2025), who analyze frictions and signaling mechanisms in matching markets. Our paper extends this line of inquiry by examining how generative AI reshapes the informativeness of signals on these platforms. More broadly, our contribution also connects to research on AI’s impact on labor markets: Agrawal et al. (2019), Noy and Zhang (2023), Brynjolfsson et al. (2025), and Acemoglu (2025) show how AI affects productivity and innovation, while we highlight an alternative channel through labor market signaling.
@@
Recent research highlights the disruptive effects of generative AI on worker behavior and employer decision-making. Dillon et al. (2025) show that generative AI reshapes productivity, information processing, and work patterns among knowledge workers. Studies closer to our context, such as Cao and Zhu (2025a), Cao and Zhu (2025b), and Chen et al. (2025), demonstrate how AI adoption changes labor market dynamics in distinct ways: Cao and Zhu (2025a) highlight fairness concerns in algorithmic decision-making and show that AI tools can exacerbate existing inequalities if not carefully designed; Cao and Zhu (2025b) document how the availability of AI writing assistance reshapes workers’ incentives to invest in communication skills and alters the distribution of opportunities across applicants; and Chen et al. (2025) argue that widespread AI use may trap workers in a “mediocrity equilibrium,” reducing incentives for effort and originality. Together, these studies underscore that the effects of AI adoption extend beyond productivity, fundamentally altering skill demand, redistributing opportunities, and influencing worker motivation. In contrast, our findings show that in the context of online labor markets, generative AI directly reshapes workers’ ability to signal their suitability to employers through cover letters, a mechanism that also connects to the broader literature on human–AI interaction. In economics more broadly, Agarwal et al. (2023) study collaboration between AI and radiologists in medical diagnosis, while in computer science, Hosanagar and Ahn (2024) and Chakrabarty et al. (2025) demonstrate how human editing can improve the quality of LLM-generated texts. Together, these strands situate our contribution within both the literature on generative AI in labor markets and the broader study of human–AI collaboration.
@@
Section 2 describes the market and the rollout of the AI tool. Section 3 presents the data, our measure of cover letter tailoring, and tool adoption patterns. Section 4 presents a simple theoretical framework that makes predictions about the effects of the AI tool. Section 5 studies the individual-level impact of generative AI on cover letters and hiring outcomes. Section 6 turns to market-level consequences, including those on cover letters as a screening device. Section 8 concludes.
@@
The AI Bid Writer tool was not the only generative AI tool that workers had access to in the sample period. One alternative is GPT-4, a general-purpose AI tool that OpenAI released on March 14, 2023, just over a month before the launch of the AI Bid Writer. Compared to its predecessor GPT-3.5 (released prior to our sample period in November 30, 2022), GPT-4 exhibits substantially enhanced reasoning. GPT‑4’s gains in logical coherence, nuanced query interpretation, and adherence to complex, long‑form instructions are well-documented in benchmark evaluations (OpenAI (2023)). Beyond raw reasoning, GPT-4 also shows advances on writing tasks, including cover-letter generation, and exhibits emergent theory-of-mind behavior that may support more audience-aware drafting (Deveci et al. (2023), Sikander et al. (2023), Kosinski (2023)).
@@
By constructing this textual alignment measure, we obtain a quantitative metric of how closely a freelancer’s cover letter mirrors the employer’s job posting, allowing for an analysis of how cover letter quality evolves over time and in response to AI. This measure, while simple, has several advantages. First, it is highly computationally inexpensive. Second, it is largely insensitive to word count and is normalized by document length—hence, variation in the word counts of job descriptions or cover letters do not introduce bias. Third, it is insensitive to word order or sentence structure, and hence explicitly focuses on content similarity while ignoring stylistic variations. [^8] Wiles et al. (2023) study the impact of an AI tool that primarily corrects spelling and grammar errors in a similar context.
@@
Another potential concern is that the rollout of the AI cover-letter tool might have coincided with changes in job postings. Figure 3 shows that the weekly number of new jobs in our sample remained stable around the launch, with a decline only later in the summer months—a pattern consistent with the drop in membership purchases. [^9] This seasonal slowdown is a well-documented phenomenon in freelance markets. Numerous specialized media reports note this cyclical decline; see, for example, Robinson (2024), Goodwin (2025), and Vozza (2016).
@@
where $\nu_{ij}\sim\mathcal{N}(0,\sigma^{2})$ is an i.i.d. shock and $q_{i}\sim\mathcal{N}(\mu_{0},\tau^{2})$ represents worker productivity. We assume that $q_{i}$, $\nu_{ij}$ and $a_{i}$ are independent, i.e. AI access is random and unrelated to productivity [^10] This formulation follows the cover-letter production function in Cowgill et al. (2024), with the modification that AI access at rollout is incomplete: only a subset of workers receive access, and for those workers the tool’s effect is an additive shift of magnitude $A$., and that $(q_{i},\quad\rho_{i},\quad\nu_{ij})$ are independent across workers. This model assumes full compliance: workers with access always use the tool. But one could easily adapt the model to one of random partial compliance.
@@
We consider the $51\%$ increase in the probability of receiving a callback to be substantial, and comparable in magnitude to effects documented in prior work on the determinants of callback rates. For instance, Bertrand and Mullainathan (2004) find that changing a job applicant’s name from an African American–sounding name to a white-sounding name raises the callback rate by about $50\%$. Our results suggest that generic cover letters represent an important barrier to receiving callbacks. Tools that help workers quickly customize their cover letters enable them to better signal interest in the project and convey alignment with the role, thereby increasing their chances of passing the initial screening stage and being invited to interview. At least at the individual level, workers clearly benefit from the AI writing tool because it increases their chances of advancing toward employment.
@@
To assess whether employers indeed substituted toward this signal, we re-estimate our baseline specification replacing cover letter tailoring with bid score rank percentile as the key regressor. The results are presented in Table 6. We find that, following the rollout of the AI tool, the predictive power of bid score rank percentile for interviews increases substantially. Specifically, the correlation between rank percentile and the probability of being interviewed rises by roughly 5% relative to the pre-period. By contrast, we detect no corresponding increase in the predictive power of rank percentile for ultimate hiring decisions. This divergence suggests that employers increasingly rely on platform-determined ranks when forming shortlists of candidates to engage with, but that final hiring outcomes continue to depend on additional factors beyond platform-calculated rankings.

### v2

@@ Abstract
We study the impact of generative AI on labor market signaling using the introduction of an AI-powered cover letter writing tool on a large online labor platform. Our data track both access to the tool and usage at the application level. Difference-in-differences estimates show that access to the tool increased textual alignment between cover letters and job posts and raised callback rates. Time spent editing AI-generated cover letter drafts is positively correlated with hiring success. After the tool’s introduction, the correlation between cover letters’ textual alignment and callbacks fell by 51%, consistent with what theory predicts if the AI technology reduces the signal content of cover letters. In response, employers shifted toward alternative signals, including workers’ prior work histories.
@@ Introduction
Job seekers and college applicants increasingly turn to generative artificial intelligence (AI) tools like ChatGPT to edit résumés, craft cover letters, and prepare for interviews. [^3] This trend has received widespread media attention. Recent headlines include “AI is supposed to make applying to jobs easier—but it might be creating another problem” (NBC News, November 17, 2024) and “AI Is Taking Over College Admissions” (The Nation, October 11, 2024). In 2025, LinkedIn reported a 45 percent year-over-year increase in job applications, with generative AI widely believed to be a key driver of the surge (“Employers Are Buried in A.I.-Generated Résumés,” The New York Times, June 21, 2025). Despite the transformative potential of this technology for processes central to economic opportunities and social mobility, credible empirical evidence on how AI affects labor market signaling remains scarce. [^4] Notable exceptions include cowgill2024does, who provide evidence on applicants’ use of generative AI in a lab setting, and wiles2025generative, who study the effect of employers’ use of generative AI tools to write job posts on an similar online labor platform. The main obstacle is observability: outside of controlled lab settings, just as employers struggle to identify AI-assisted applications, so too do researchers. Current detection tools are error-prone and unreliable (weber2023testing), making it difficult to assess AI assistance’s real-world impact on hiring.
@@
To measure how closely cover letters are tailored to job descriptions, we apply Term Frequency–Inverse Document Frequency (TF-IDF), a standard Natural Language Processing technique that captures the extent of keyword overlap between two texts. Using data from before the introduction of the AI writing tool, we show that our TF-IDF–based measure of cover-letter tailoring similarly predicts callbacks and awards—both across and within workers—providing complementary evidence that employers value cover letters that engage with job content.
@@
We find that access to the generative AI writing tool increased cover-letter tailoring by 0.16 standard deviations, while actual usage raised tailoring by 1.36 standard deviations. Applying the same design to callbacks as the outcome, we find that access to the generative AI tool increased the probability of receiving a callback by 0.43 percentage points, and usage raised it by 3.56 percentage points. The latter represents a 51% increase relative to the pre-rollout average callback rate of 7.02%. However, the results for callbacks are significant at the $10\%$ level but not at the $5\%$ level. A dynamic effect specification shows that the effect on callbacks tapered off after two months. These results suggest that generic or poorly matched cover letters present a significant barrier to receiving callbacks. In contrast, the estimates on actual awards are too imprecise to draw definitive conclusions.
Our second finding is that AI substitutes for, rather than complements, workers’ pre-AI cover letter tailoring skills. We compute workers’ average tailoring scores prior to the tool’s introduction and study the heterogeneous effects of access to the tool by pre-existing writing ability. We find that workers who previously wrote more tailored cover letters experienced smaller gains in cover letter tailoring—indeed, the best writers gained half as much as the weakest ones. By enabling less skilled writers to produce more tailored cover letters, AI narrows the gap between workers with different initial abilities.
@@
galdin2025making offer complementary evidence on the effects of generative AI on signaling in the same labor market. They use an LLM-based measure of cover letter quality and find, consistent with our results, that its correlation with application outcomes declined after AI tools became available. While we provide difference-in-differences evidence on the platform policy and evidence on human–AI interaction, they estimate an equilibrium model and perform counterfactual simulations.
@@
Another closely related paper is cowgill2024does, who examine the impact of generative AI on job applications and startup pitches through a survey experiment. They find that generative AI reduces employers’ and investors’ screening accuracy. Our study explores similar questions in an online labor market, examining the effect of AI on actual hiring decisions. An additional important difference between the settings is that cowgill2024does informed employers whether applicants used generative AI, whereas, in many real-world contexts—including ours—employers typically cannot discern whether an applicant received assistance from AI. As a result, our findings are not directly comparable.
@@
In a similar labor market, wiles2023algorithmic study the effect of a non-generative algorithmic writing assistance on employment through an experiment. The intervention they study focuses mostly on correcting spelling and grammar errors. They show that the tool improved treated workers’ employment outcomes without affecting the employers’ satisfaction. While we also investigate the role of AI—although of a different nature—in workers’ employment outcomes, we are additionally able to speak to the market-level outcomes associated with the introduction of generative AI. Another related paper is wiles2025generative, who study generative AI assistance to employers for crafting job posts in a similar context and find that while the tool increased the number of job posts, it did not increase the number of matches. Our work is complementary to theirs as both sides of the labor market—firms and workers—increasingly adopt generative AI tools in matching.
@@
Also related is work that examines the functioning of online labor markets and their broader economic implications. horton2010online, chen2016are, and stanton2020gig document how these markets operate, the behavioral responses of workers to wage changes, and the significance of platform-mediated employment. stanton2025who provide evidence on the distributional consequences of gig platforms, showing how access to online markets affects both workers and firms, and highlighting heterogeneity in who ultimately benefits from platform participation. This focus on heterogeneous impacts complements theoretical and empirical work such as horton2024sorting and allman2025signaling, who analyze frictions and signaling mechanisms in matching markets. Our paper extends this line of inquiry by examining how generative AI reshapes the informativeness of signals on these platforms. More broadly, our contribution also connects to research on AI’s impact on labor markets: agrawal2019artificial, noy2023experimental, brynjolfsson2025generative, and acemoglu2025simple show how AI affects productivity and innovation, while we highlight an alternative channel through labor market signaling. Also related is brinatti2023international, who use data from a global online labor platform to study wage determination in remote work.
@@
Recent research highlights the disruptive effects of generative AI on worker behavior and employer decision-making. dillon2025shifting show that generative AI reshapes productivity, information processing, and work patterns among knowledge workers. Studies closer to our context, such as cao2025fairness, cao2025aiwrites, and chen2025mediocrity, demonstrate how AI adoption changes labor market dynamics in distinct ways: cao2025fairness highlight fairness concerns in algorithmic decision-making and show that AI tools can exacerbate existing inequalities if not carefully designed; cao2025aiwrites document how the availability of AI writing assistance reshapes workers’ incentives to invest in communication skills and alters the distribution of opportunities across applicants; and chen2025mediocrity argue that widespread AI use may trap workers in a “mediocrity equilibrium,” reducing incentives for effort and originality. Together, these studies underscore that the effects of AI adoption extend beyond productivity, fundamentally altering skill demand, redistributing opportunities, and influencing worker motivation. In contrast, our findings show that in the context of online labor markets, generative AI directly reshapes workers’ ability to signal their suitability to employers through cover letters, a mechanism that also connects to the broader literature on human–AI interaction. In economics more broadly, agarwal2023combining study collaboration between AI and radiologists in medical diagnosis, while in computer science, hosanagar2024designing and chakrabarty2025can demonstrate how human editing can improve the quality of LLM-generated texts. Together, these strands situate our contribution within both the literature on generative AI in labor markets and the broader study of human–AI collaboration.
In developing our model, we build most directly on the framework of cowgill2024does, who study how the introduction of generative AI affects the informativeness of job applications. Their analysis shows that when AI assistance becomes available, application materials become less revealing of underlying worker quality, reshaping how employers interpret signals in hiring. Our model is also similar to that of frankel2019muddled in which quality can be interpreted as a natural action and access to AI gaming ability. Together, these papers highlight the central insight that new technologies or multidimensional private information can weaken the mapping between observed signals and underlying productivity, and our model applies this perspective to the case of AI-assisted cover letters.
@@
Section 2 describes the market and the rollout of the AI tool. Section 3 presents the data, our measure of cover letter tailoring, and tool adoption patterns. Section 4 presents a simple theoretical framework that makes predictions about the effects of the AI tool. Section 5 studies the individual-level impact of generative AI on cover letters and hiring outcomes. Section 6 turns to market-level consequences, including those on cover letters as a screening device. Section 7 examines the human-AI interaction. Section 8 concludes.
@@
The AI Bid Writer tool was not the only generative AI tool that workers had access to in the sample period. One alternative is GPT-4, a general-purpose AI tool that OpenAI released on March 14, 2023, just over a month before the launch of the AI Bid Writer. Compared to its predecessor GPT-3.5 (released prior to our sample period in November 30, 2022), GPT-4 exhibits substantially enhanced reasoning. GPT‑4’s gains in logical coherence, nuanced query interpretation, and adherence to complex, long‑form instructions are well-documented in benchmark evaluations (OpenAI2023GPT4). Beyond raw reasoning, GPT-4 also shows advances on writing tasks, including cover-letter generation, and exhibits emergent theory-of-mind behavior that may support more audience-aware drafting (Deveci2023CoverLetters, Sikander2023Intro, Kosinski2023ToM).
@@
By constructing this textual alignment measure, we obtain a quantitative metric of how closely a freelancer’s cover letter mirrors the employer’s job posting, allowing for an analysis of how cover letter quality evolves over time and in response to AI. This measure, while simple, has several advantages. First, it is highly computationally inexpensive. Second, it is largely insensitive to word count and is normalized by document length—hence, variation in the word counts of job descriptions or cover letters do not introduce bias. Third, it is insensitive to word order or sentence structure, and hence explicitly focuses on content similarity while ignoring stylistic variations.777wiles2023algorithmic study the impact of an AI tool that primarily corrects spelling and grammar errors in a similar context.
@@
Another potential concern is that the rollout of the AI cover-letter tool might have coincided with changes in job postings. Figure 3 shows that the weekly number of new jobs in our sample remained stable around the launch, with a decline only later in the summer months—a pattern consistent with the drop in membership purchases. [^8] This seasonal slowdown is a well-documented phenomenon in freelance markets. Numerous specialized media reports note this cyclical decline; see, for example, robinson2024summer, goodwin2025survive, and vozza2016summer.
@@
where $\nu_{ij}\sim\mathcal{N}(0,\sigma^{2})$ is an i.i.d. shock and $q_{i}\sim\mathcal{N}(\mu_{0},\tau^{2})$ represents worker productivity. We assume that $q_{i}$, $\nu_{ij}$ and $a_{i}$ are independent, i.e. AI access is random and unrelated to productivity [^9] This formulation follows the cover-letter production function in cowgill2024does, with the modification that AI access at rollout is incomplete: only a subset of workers receive access, and for those workers the tool’s effect is an additive shift of magnitude $A$., and that $(q_{i},\quad\rho_{i},\quad\nu_{ij})$ are independent across workers. This model assumes full compliance: workers with access always use the tool. But one could easily adapt the model to one of random partial compliance.
@@
We consider the $51\%$ increase in the probability of receiving a callback to be substantial, and comparable in magnitude to effects documented in prior work on the determinants of callback rates. For instance, bertrand2004emily find that changing a job applicant’s name from an African American–sounding name to a white-sounding name raises the callback rate by about $50\%$. Our results suggest that generic cover letters represent an important barrier to receiving callbacks. Tools that help workers quickly customize their cover letters enable them to better signal interest in the project and convey alignment with the role, thereby increasing their chances of passing the initial screening stage and being invited to interview. At least at the individual level, workers clearly benefit from the AI writing tool because it increases their chances of advancing toward employment.
@@
To assess whether employers indeed substituted toward this signal, we re-estimate our baseline specification replacing cover letter tailoring with bid score rank percentile as the key regressor. The results are presented in Table 6. We find that, following the rollout of the AI tool, the predictive power of bid score rank percentile for interviews increases substantially. Specifically, the correlation between rank percentile and the probability of being interviewed rises by roughly 5% percent relative to the pre-period. By contrast, we detect no corresponding increase in the predictive power of rank percentile for ultimate hiring decisions. This divergence suggests that employers increasingly rely on platform-determined ranks when forming shortlists of candidates to engage with, but that final hiring outcomes continue to depend on additional factors beyond platform-calculated rankings.